# InPK: Infusing Prior Knowledge into Prompt for Vision-Language Models

### v1

@@ Abstract
Prompt tuning has become a popular strategy for adapting Vision-Language Models (VLMs) to zero/few-shot visual recognition tasks. Some prompting techniques introduce prior knowledge due to its richness, but when learnable tokens are randomly initialized and disconnected from prior knowledge, they tend to overfit on seen classes and struggle with domain shifts for unseen ones. To address this issue, we propose the InPK model, which infuses class-specific prior knowledge into the learnable tokens during initialization, thus enabling the model to explicitly focus on class-relevant information. Furthermore, to mitigate the weakening of class information by multi-layer encoders, we continuously reinforce the interaction between learnable tokens and prior knowledge across multiple feature levels. This progressive interaction allows the learnable tokens to better capture the fine-grained differences and universal visual concepts within prior knowledge, enabling the model to extract more discriminative and generalized text features. Even for unseen classes, the learned interaction allows the model to capture their common representations and infer their appropriate positions within the existing semantic structure. Moreover, we introduce a learnable text-to-vision projection layer to accommodate the text adjustments, ensuring better alignment of visual-text semantics. Extensive experiments on 11 recognition datasets show that InPK significantly outperforms state-of-the-art methods in multiple zero/few-shot image classification tasks.
1 Introduction
@@
In recent years, Vision-Language Models (VLMs) pre-trained on large-scale image-text datasets, such as CLIP [25], have shown strong generalization across downstream tasks. To improve model adaptation to specific tasks or datasets, particularly in zero/few-shot settings, the community has proposed many solutions [8, 44, 50, 49, 13, 14]. Among these, prompt tuning [50, 49, 13, 14] has emerged as a popular strategy for adapting VLMs to various visual recognition tasks in zero/few-shot scenarios.
@@
Figure 1: (a) Existing prior knowledge-based prompting techniques initialize learnable tokens randomly, leaving them disconnected from class-specific prior knowledge (illustrated in top). Our approach infuses prior knowledge into learnable tokens before feeding them into the encoder layers each time, providing a prior knowledge-based initialization that explicitly emphasizes class-relevant information (illustrated in bottom). (b) t-SNE visualization of the feature manifolds for our method with and without prior knowledge infusion (pki) on the EuroSAT dataset. Our method with pki shows tighter intra-class distances, indicating enhanced consistency, and larger inter-class distances, reflecting improved class discrimination. (c) In the base-to-base/base-to-novel setting, our method significantly outperforms the state-of-the-art methods regarding average results on 11 recognition datasets.
@@
Prompt tuning techniques have gained attention for their ability to improve VLMs’ performance on downstream tasks by leveraging textual prompts. For example, CoOp [50] and CoCoOp [49] replace fixed prompts with learnable tokens, achieving substantial improvements through fine-tuning on a limited number of samples. Recent works [11, 14, 13] extend this concept to multimodal prompt tuning, fine-tuning both text and visual branches to improve performance on both base (seen) and novel (unseen) classes. However, since learnable tokens are optimized primarily for the base classes during training, these methods often suffer from overfitting, which leads to domain shifts when encountering novel classes. To address this problem, a line of works [16, 21, 28, 46, 12] introduce prior knowledge to enhance object understanding, helping the model better match visual information with textual descriptions. While these approaches alleviate overfitting to some extent, the random initialization of learnable tokens and their disconnection from prior knowledge (Fig. 1:a (top)) still hinder the model’s ability to extract discriminative and generalized features, limiting its performance improvement on both base and novel classes.
To address the aforementioned issue, we strategically infuse the class-specific prior knowledge into learnable tokens at the initialization stage (Fig. 1:a (bottom)) to establish explicit class-relevant preferences and provide higher-quality input to subsequent layers. Furthermore, to prevent the dilution of class information across multi-layer encoders, we continuously reinforce the interaction between learnable tokens and prior knowledge at various feature levels. This progressive design enables the model to more effectively focus on class-relevant details. For base classes, this design reduces overfitting to irrelevant information during training, enhancing the model’s ability to perceive fine-grained differences and extract more discriminative text features, thus decreasing the likelihood of misclassification. For novel classes, learnable tokens acquire generalization capabilities through common visual concepts derived from prior knowledge, helping the model infer the appropriate position of novel classes within the existing semantic structure and thereby improving its generalization performance on unseen classes. Moreover, while prior knowledge-based methods often prioritize the text branch, they tend to overlook the impact of text adjustments on visual semantic alignment. To address this issue, we introduce a learnable text-to-vision projection layer to enforce stronger alignment between visual and textual semantics.
In this paper, we propose InPK, a novel approach that infuses prior knowledge into prompt for vision-language models. The t-SNE visualization results also demonstrate the effectiveness of prior knowledge infusion. (Fig. 1:b). Given the importance of high-quality prior knowledge in enhancing category recognition, and the rich world knowledge inherent in Large Language Models (LLMs), we leverage LLMs to generate comprehensive and discriminative prior knowledge using a predefined instruction template. This knowledge consists of detailed attribute terms that capture subtle class features, refining the model’s understanding of class boundaries. It is then infused into learnable tokens at multiple feature levels, with progressively strengthened interactions, enabling the model to learn more discriminative and generalized features. Moreover, since the text branch incorporates prior knowledge, we introduce a learnable text-to-vision projection layer to adapt to text adjustments, ensuring better alignment between visual and textual semantics. To address potential confusion caused by common attribute words across different categories, we introduce regularization constraints that emphasize the importance of class names, preventing the model from overfitting to shared attributes. Notably, experimental results of average accuracy on 11 popular recognition datasets show that InPK significantly outperforms state-of-the-art methods on the base and novel classes (Fig. 1:c). The main contributions of this paper include:
@@
  • We infuse prior knowledge into learnable tokens across multiple feature levels, guiding the model to extract discriminative and generalized features, therefore improving performance on both base and novel classes.
  • We introduce a learnable text-to-vision projection layer to accommodate the text adjustments, enhancing the alignment between visual and textual semantics.
@@
  • Our method conducts extensive experiments on 11 popular visual recognition datasets, demonstrating its effectiveness in multiple zero/few-shot image classification tasks.
2 Related Work
2.1 Vision-Language Models
@@
Vision-Language Models (VLMs) represent significant progress in multimodal learning [30, 31, 29] in recent years. These models are pre-trained on large-scale datasets to effectively learn joint representations from both images and text. Recent studies [25, 10, 39, 40, 18] have demonstrated VLMs’ superior performance in tasks such as zero/few-shot image recognition. Notably, the pioneering work CLIP  [25] is renowned for its simplicity and effectiveness, leveraging large-scale image-text pair training and contrastive learning techniques. Other works like ALIGN [10], FILIP [39], Florence [40], and REACT [18] further highlight VLMs’ strong open vocabulary understanding. To adapt pre-trained VLMs to specific downstream tasks, numerous task-specific methods have been proposed including segmentation [33, 45, 48], image recognition [43, 1, 37, 36, 32] and object detection [6, 41, 51, 47]. In this work, we propose a novel prompt-tuning technique that enhances VLMs’ generalization performance across various visual recognition tasks.
2.2 Prompt Learning for Vision-Language Models
@@
Fine-tuning Vision-Language Models (VLMs) for downstream tasks or datasets while preserving their original generalization capabilities poses significant challenges. Retraining the entire model is often impractical due to the vast number of parameters, and it risks overfitting, which could diminish the generalization benefits obtained during pre-training. To tackle this, prompt-tuning methods have emerged, with CoOp  [50] being a pioneering approach. CoOp adapts VLMs to downstream tasks by introducing learnable vectors in place of hand-crafted prompts. Extensions of this approach [49, 13, 42, 37, 36, 23], multimodal prompt tuning [13, 19, 3] has been applied to fine-tune both text and visual branches simultaneously. To prevent overfitting, a line of works[14, 37, 38, 52] have introduced regularization constraints to reduce the loss of general information. Other approaches, such as UNIGRAM [17] and ProMetaR [23], leverage meta-learning to enhance generalization by initializing prompts or applying regularization in the meta-learning framework. More recently, some methods [16, 21, 28, 46, 12] have introduced external knowledge to enrich textual representations and better capture complex visual semantics. Our approach goes beyond using learnable tokens by infusing them into class-specific prior knowledge to provide class-relevant preferences, resulting in improved generalization performance on downstream tasks.
3 Methodology
In this section, we first provide a brief overview of CLIP [25]. Next, we comprehensively introduce our proposed method InPK.
3.1 Review of CLIP
@@
3.2 InPK: Infusing Prior Knowledge into Prompt
@@
Based on the CLIP model, to further adapt the VLMs to specific downstream tasks while excelling in both base and novel classes, we propose the InPK model, as depicted in Fig. 2. Different from other methods based on prior knowledge [16, 21, 46, 12], which randomly initialize the learnable tokens, our approach infuses prior knowledge into learnable tokens at the initialization stage and progressively strengthens the connection between prior knowledge and tokens across multiple feature levels. Furthermore, we introduce a learnable text-to-vision projection layer to accommodate text adjustments and enforce alignment between visual and textual semantics. Finally, regularization constraint is employed with the objective of retaining general knowledge and emphasizing the role of class names.
3.2.1 Generation of Prior Knowledge
@@
Regarding prior knowledge, relying solely on class names, especially for rare or ambiguous scientific terms, limits the model’s ability to effectively associate these terms with visual information. Therefore, it is essential to incorporate attribute words related to each category alongside the class names. In our method, prior knowledge includes both class names and class-specific attributes, enriching the model’s semantic understanding and universal visual concepts brought by attributes that provide cross-class knowledge. We generate attribute words using GPT-4, leveraging its rich world knowledge and guide the process with a predefined instruction template that incorporates both the class name (fine-grained) and dataset name (coarse-grained). The template is as follows: “Please provide the discriminative features about the [fine class] for recognition among all [coarse class]”. This process yields the input $\boldsymbol{G}=\{\boldsymbol{c_{k},g_{1},\dots,g_{N}}\}$, where $\boldsymbol{c_{k}}$ denotes the class name for the $k$-th class, and $\{\boldsymbol{g_{1},\dots,g_{N}}\}$ represent $N$ attribute words. Prior knowledge is generated offline, eliminating the need for training or inference time. Samples of these attribute words are provided in the supplementary material.
3.2.2 Prior Knowledge-infused Text Encoder
@@
With the acquired class-specific prior knowledge, we avoid merely concatenating the learnable tokens with prior knowledge before feeding them into the text encoder. Instead, our approach infuses the prior knowledge into the learnable tokens, introducing explicit category preference and delivering higher-quality input to subsequent layers. Therefore, beyond the original feature $\boldsymbol{f}$ derived from hand-crafted prompts in the text branch, we introduce an additional pathway within the PKi module that incrementally models features, ultimately producing the new textual feature $\boldsymbol{\widetilde{f}}$. Specifically, the input tokens for the PKi module are structured as $\{\boldsymbol{P_{t},\widetilde{E}}\}$, where $\boldsymbol{P_{t}}=\{\boldsymbol{p_{t}^{1},p_{t}^{2},\dots,p_{t}^{M}}\}$ consists of $M$ learnable tokens, and $\boldsymbol{\widetilde{E}}=\{\boldsymbol{\widetilde{e}_{1},\widetilde{e}_{2},\ldots,\widetilde{e}_{n}}\}$ represents the word embeddings of $\boldsymbol{G}$. We then introduce an attribute-aware attention mechanism that fuses class-specific knowledge with the learnable tokens. In this mechanism, the learnable tokens $\boldsymbol{P_{t}}$ serve as the query $Q$, while the class-specific prior knowledge embeddings $\boldsymbol{\widetilde{E}}$ as both the key $K$ and the value $V$, resulting in a robust token embedding $\boldsymbol{\widehat{P_{t}}}$ that incorporates prior knowledge:
@@
$\boldsymbol{\widehat{P_{t}}}=\textit{FFN}(\textit{LN}(\boldsymbol{A}))+\boldsymbol{A},$ (5)
where MultiHead () and FFN () follow the standard Transformer configurations [27], representing multi-head attention and feed-forward neural network, respectively. Layer normalization (LN) is applied before each block. Furthermore, to counteract the potential dilution of category information caused by multi-layer encoders, we incorporate attribute-aware attention into each transformer layer of the PKi up to a specific depth $J$, allowing for step-by-step feature modeling. Specifically, the PKi itself comprises $L$ layers of frozen transformers, with all parameters pre-trained. The fused learnable token $\boldsymbol{\widehat{P}_{t}^{i}}$ is concatenated with the $\boldsymbol{E^{i}}$, and the resulting tokens are then passed to the $(i+1)^{\mathrm{th}}$ transformer layer:
@@
3.2.3 Alignment of Text and Visual Branches

### v2

@@ Abstract
Prompt tuning has emerged as a prevalent approach for adapting VLMs to downstream visual recognition tasks in zero/few-shot settings. While some prompting techniques leverage prior knowledge for its richness, the random initialization and disconnection of learnable tokens from this knowledge often lead to overfitting to seen classes and poor generalization across unseen domains. To address this issue, we propose the InPK model, which infuses class-specific prior knowledge into the learnable tokens during initialization, thus enabling the model to explicitly focus on class-relevant information. Furthermore, to mitigate the weakening of class information by multi-layer encoders, we continuously reinforce the interaction between learnable tokens and prior knowledge across multiple feature levels. This progressive interaction allows the learnable tokens to better capture the fine-grained differences and universal visual concepts within prior knowledge, enabling the model to extract more discriminative and generalized features. Even for unseen classes, the learned interaction allows the model to capture their common representations and infer their appropriate positions within the existing semantic structure. Moreover, we introduce a learnable text-to-vision projection layer to accommodate the text adjustments, ensuring better alignment of visual-text semantics. Extensive experiments on 11 recognition datasets show that InPK significantly outperforms state-of-the-art methods in multiple zero/few-shot image classification tasks.
Index Terms:
Prompt tuning, vision-language model, few-shot classification.
I Introduction
@@
In recent years, Vision-Language Models (VLMs) pre-trained on large-scale image-text datasets, such as CLIP [1], have demonstrated strong generalization capabilities across downstream tasks. Typically, VLMs employ contrastive loss to optimize both image and text encoders, aligning features from both modalities within a shared embedding space. To further enhance model adaptation to specific tasks or datasets, particularly in zero/few-shot settings, the community has proposed numerous solutions [2, 3, 4, 5, 6, 7]. Among these, prompt tuning [4, 5, 6, 7] has emerged as a popular strategy for adapting VLMs to various visual recognition tasks in zero/few-shot scenarios.
@@
Figure 1: (a) Existing prior knowledge-based prompting techniques initialize learnable tokens randomly, resulting in task-agnostic token representations that prolong the optimization trajectory and increase susceptibility to local optima (illustrated in top). Our approach infuses prior knowledge into learnable tokens before feeding them into the encoder layers each time, providing a prior knowledge-based initialization that explicitly emphasizes class-relevant information (illustrated in bottom). (b) t-SNE visualization of the feature manifolds for our method with and without prior knowledge infusion (pkin) on the EuroSAT dataset. Our method with pkin shows tighter intra-class distances, indicating enhanced consistency, and larger inter-class distances,n reflecting improved class discrimination. (c) In the base-to-base/base-to-novel setting, our method significantly outperforms the state-of-the-art methods regarding average results on 11 recognition datasets.
@@
Prompt tuning techniques have gradually gained favor in adapting Vision and Language Models (VLMs) to downstream tasks without the need to fully retrain the original models. For example, CoOp [4] and CoCoOp [5] replace fixed prompts with learnable tokens, achieving substantial improvements through fine-tuning on a limited number of samples. Recent works [8, 7, 6, 9] extend this concept to multimodal prompt tuning, fine-tuning both text and visual branches to improve performance on both base (seen) and novel (unseen) classes. However, since learnable tokens are primarily optimized for base classes during training, prompt-based models often overfit to task-specific data distributions.
Recently, a series of studies [10, 11, 12, 13, 14] introduce prior knowledge to enhance object understanding and build more coherent text-visual mappings in the semantic space. While these methods mitigate overfitting to some extent, the random initialization of learnable tokens and their disconnection from prior knowledge (see Figure 1: a (top)) remain limiting factors. First, random initialization may cause prompt tokens to fall into semantically irrelevant regions, leading to an optimization starting point misaligned with the target semantic space and increasing the risk of converging to suboptimal solutions. Second, as training progresses, the static representation of prior knowledge and dynamically optimized learnable tokens tend to exhibit representation drift. This shift becomes particularly pronounced in few-shot scenarios, where the model excessively relies on learnable tokens to fit base class features, failing to effectively leverage prior knowledge for representation adjustment, ultimately degrading cross-class generalization. Furthermore, although learnable tokens and prior knowledge may interact implicitly in subsequent network layers (e.g., Transformer self-attention), the lack of explicit constraints makes it difficult to suppress semantic noise and enhance semantic relevance in the high-dimensional embedding space. Consequently, these limitations still hinder the model’s ability to extract discriminative and generalized features, constraining its performance on both base and novel classes.
To address the aforementioned issues, we propose InPK (Infusing Prior Knowledge into Prompt), which enhances the generalization capability of vision-language models through a hierarchical prior knowledge infusion mechanism. During initialization, we inject class-specific prior knowledge generated by a large language model (LLM) into the learnable tokens, establishing class-related semantic preferences and constraining the tokens within a semantically meaningful subspace (see Figure 1: a (bottom)). For base classes, this design encourages the model to focus on essential features aligned with prior knowledge, enhancing sensitivity to subtle distinctions. For novel classes, the learnable tokens create cross-class semantic bridges through shared visual concepts, allowing the feature representations of novel classes to adaptively integrate into the existing semantic space, thereby improving generalization to unseen classes [15]. Visualization shows that compared with the method of directly concatenating prior knowledge and learnable tokens, the feature space produced by InPK has tighter intra-class clustering and greater inter-class separation (see Figure 1: b), verifying the effectiveness of our prior knowledge fusion strategy (see Figure 1: c).
Furthermore, to counteract the gradual dilution of class information across multi-layer encoders, we continuously reinforce the interaction between the learnable tokens and prior knowledge at various feature levels, progressively capturing fine-grained vision-semantic associations. For more precise semantic alignment between textual and visual modalities, we introduce a learnable cross-modal projection layer. Moreover, InPK improves model robustness via a regularization mechanism by incorporating a Class-Name Saliency Regularization term into the loss function, encouraging the model to focus on core class semantics while mitigating over-reliance on shared attributes. Notably, experimental results of average accuracy on 11 popular recognition datasets show that InPK significantly outperforms state-of-the-art methods on the base and novel classes.
The main contributions of this paper include:
@@
  • We infuse prior knowledge into learnable tokens at multiple feature levels, guiding the model to extract more discriminative and generalizable features, thereby enhancing recognition performance on both base and novel classes.
  • We introduce a learnable text-to-vision projection layer to accommodate the text prompt tuning process, thereby enhancing vision-text semantic alignment and improving cross-modal consistency.
@@
  • We conduct extensive experiments on 11 popular vision recognition datasets, and the results demonstrate that InPK achieves significant performance improvements across multiple zero/few-shot image classification tasks.
II Related Work
II-A Vision-Language Models
@@
Vision-Language Models (VLMs) represent significant progress in multimodal learning [16, 17, 18] in recent years. These models are pre-trained on large-scale datasets to effectively learn joint representations from both images and text. Recent studies [1, 19, 20, 21, 22] have demonstrated VLMs’ superior performance in tasks such as zero/few-shot image recognition. Notably, the pioneering work CLIP  [1] is renowned for its simplicity and effectiveness, leveraging large-scale image-text pair training and contrastive learning techniques. Other works like ALIGN [19], FILIP [20], Florence [21], and REACT [22] further highlight VLMs’ strong open vocabulary understanding. To adapt pre-trained VLMs to specific downstream tasks, numerous task-specific methods have been proposed including segmentation [23, 24, 25, 26], image recognition [27, 28, 29, 30, 31, 32] and object detection [33, 34, 35, 36, 37]. In this work, we propose a novel prompt-tuning technique that enhances VLMs’ generalization performance across various visual recognition tasks.
II-B Prompt Learning for Vision-Language Models
@@
Fine-tuning Vision-Language Models (VLMs) for downstream tasks or datasets while preserving their original generalization capabilities poses significant challenges. Retraining the entire model is often impractical due to the vast number of parameters, and it risks overfitting, which could diminish the generalization benefits obtained during pre-training. To tackle this, prompt-tuning methods have emerged, with CoOp  [4] being a pioneering approach. CoOp adapts VLMs to downstream tasks by introducing learnable vectors in place of hand-crafted prompts. Extensions of this approach, multimodal prompt tuning [6, 38, 39, 40] has been applied to fine-tune both text and visual branches simultaneously. To prevent overfitting, a line of works[7, 29, 41, 42] have introduced regularization constraints to reduce the loss of general information. Other approaches, such as UNIGRAM [43] and ProMetaR [44], leverage meta-learning to enhance generalization by initializing prompts or applying regularization in the meta-learning framework. More recently, some methods [10, 11, 12, 15, 13] have introduced external knowledge to enrich textual representations and better capture complex visual semantics. Our approach goes beyond using learnable tokens by infusing them into class-specific prior knowledge to provide class-relevant preferences, resulting in improved generalization performance on downstream tasks.
III Methodology
In this section, we first provide a brief overview of CLIP [1], followed by a comprehensive introduction to our proposed method, InPK, including the generation of prior knowledge, the prior knowledge-infused text encoder, the alignment of text and visual branches, and the training objective.
III-A Review of CLIP
@@
III-B InPK: Infusing Prior Knowledge into Prompt
@@
Based on the CLIP model, to further adapt the VLMs to specific downstream tasks while excelling in both base and novel classes, we propose the InPK model, as depicted in Figure  2. Different from other methods based on prior knowledge [10, 12, 14], which randomly initialize the learnable tokens, our approach infuses prior knowledge into learnable tokens at the initialization stage and progressively strengthens the connection between prior knowledge and tokens across multiple feature levels. Furthermore, we introduce a learnable text-to-vision projection layer to accommodate text adjustments and enforce alignment between visual and textual semantics. Finally, regularization constraint is employed with the objective of retaining general knowledge and emphasizing the role of class names.
III-B1 Generation of Prior Knowledge
@@
Regarding prior knowledge, relying solely on class names, especially for rare or ambiguous scientific terms, limits the model’s ability to effectively associate these terms with visual information. Therefore, it is essential to incorporate attribute words related to each category alongside the class names [13]. In our method, prior knowledge includes both class names and class-specific attributes, enriching the model’s semantic understanding and universal visual concepts brought by attributes that provide cross-class knowledge. We generate attribute words using GPT-4, leveraging its rich world knowledge, and guide the process with a predefined instruction template that incorporates both the class name and dataset name. The template is as follows: “Please provide the discriminative features about the [class name] for recognition”. This process yields the input $\boldsymbol{G}=\{\boldsymbol{c_{k},g_{1},\dots,g_{N}}\}$, where $\boldsymbol{c_{k}}$ denotes the class name for the $k$-th class, and $\{\boldsymbol{g_{1},\dots,g_{N}}\}$ represent $N$ attribute words. Prior knowledge is generated offline, eliminating the need for training or inference time.
III-B2 Prior Knowledge-infused Text Encoder
@@
With the acquired class-specific prior knowledge, we avoid merely concatenating the learnable tokens with prior knowledge before feeding them into the text encoder. Instead, our approach infuses prior knowledge into the learnable tokens, introducing explicit category preference and delivering higher-quality input to subsequent layers. Therefore, beyond the original feature $\boldsymbol{f}$ derived from hand-crafted prompts in the text branch, we introduce an additional pathway within the Prior Knowledge-infused text encoder(PKi) that incrementally models features, ultimately producing the new textual feature $\boldsymbol{\widetilde{f}}$. Specifically, the input tokens for the PKi module are structured as $\{\boldsymbol{P_{t},\widetilde{E}}\}$, where $\boldsymbol{P_{t}}=\{\boldsymbol{p_{t}^{1},p_{t}^{2},\dots,p_{t}^{M}}\}$ consists of $M$ learnable tokens, and $\boldsymbol{\widetilde{E}}=\{\boldsymbol{\widetilde{e}_{1},\widetilde{e}_{2},\ldots,\widetilde{e}_{n}}\}$ represents the word embeddings of $\boldsymbol{G}$. We then introduce an attribute-aware attention mechanism $\mathcal{T}$ that fuses class-specific knowledge with the learnable tokens. In this mechanism, the learnable tokens $\boldsymbol{P_{t}}$ serve as the query $Q$, while the class-specific prior knowledge embeddings $\boldsymbol{\widetilde{E}}$ as both the key $K$ and the value $V$, resulting in a robust token embedding $\boldsymbol{\widehat{P_{t}}}$ that incorporates prior knowledge:
$\boldsymbol{\widehat{P_{t}}}=\mathcal{T}(\boldsymbol{P_{t}},\boldsymbol{\widetilde{E}})=\textit{FFN}(\textit{LN}(\boldsymbol{A}))+\boldsymbol{A},$ (4)
@@
where MultiHead () and FFN () follow the standard Transformer configurations [45], representing multi-head attention and feed-forward neural network, respectively. Layer normalization (LN) is applied before each block.
Furthermore, to counteract the potential dilution of category information caused by multi-layer encoders, we incorporate attribute-aware attention into each transformer layer of the PKi up to a specific depth $J$, allowing for step-by-step feature modeling. Specifically, the PKi itself comprises $L$ layers of frozen transformers, with all parameters pre-trained. The fused learnable token $\boldsymbol{\widehat{P}_{t}^{i-1}}$ is concatenated with the $\boldsymbol{\widetilde{E}^{i-1}}$, and the resulting tokens are then passed to the $i^{\mathrm{th}}$ transformer layer:
$\boldsymbol{\widehat{P}_{t}^{i-1}}=\mathcal{T}(\boldsymbol{P_{t}^{i-1}},\boldsymbol{\widetilde{E}^{i-1}}),$ (6)
@@
III-B3 Alignment of Text and Visual Branches